If you earn a real rate of return of 2% and inflation is 4%, what is...

  1. If you earn a real rate of return of 2% and inflation is 4%, what is your nominal rate of return? (to 2 decimal places)
Answer #1

(1+ nominal rate) = (1 + real rate)*(1+inflation rate)

(1+ nominal rate) = (1 + 0.02)*(1+0.04) = 1.0608

(1 + nominal rate) = 1.0608

nominal rate = 1.0608-1 =0.0608 = 6.08%

Answer : 6.08% (THUMBS UP PLEASE)